What is the molar mass of glucose if a solution containing 18.00 g of glucose in 150 g of water boils at 100.34 degrees centigrade?


1
Expert's answer
2021-02-12T05:03:26-0500

b = "\frac{\Delta T_b}{K_b}"

b= molarity of solution

"\Delta T_b" = 100.34 - 100 = 0.34 ⁰C

Kb = ebullioscopic constant = 0.512 ⁰C Kg/mol

b = "\frac{0.34}{0.512}" = 0.664 mol/Kg

This means that you have 0.664 moles of glucose for every 1 kg of water. The number of moles you'll have in 150 g of water will be

150 g × "\frac{0.664 mol}{1000g}" = 0.0996 moles


0.0996 moles × "\frac{molar mass of glucose} {1 mole }" = 18.9 g


molar mass of glucose = 189.75 g